What volume of 0.355 M HCl is needed to make 200. mL of 0.150 M HCl ?
Answer #1

Answer

M1 = 0.355M

V1=?

M2= 0.150M

V2 = 200 mL

Using the molar equation

M1 ×V1 = M2×V2

0.355M ×V1=0.150 M×200mL

V1= 84.51mL